# AdaDetectGPT: Adaptive Detection of LLM-Generated Text with Statistical Guarantees

### v1

@@ Abstract
We study the problem of determining whether a piece of text has been authored by a human or by a large language model (LLM). Existing state of the art logits-based detectors make use of statistics derived from the log-probability of the observed text evaluated using the distribution function of a given source LLM. However, relying solely on log probabilities can be sub-optimal. In response, we introduce AdaDetectGPT – a novel classifier that adaptively learns a witness function from training data to enhance the performance of logits-based detectors. We provide statistical guarantees on its true positive rate, false positive rate, true negative rate and false negative rate. Extensive numerical studies show AdaDetectGPT nearly uniformly improves the state-of-the-art method in various combination of datasets and LLMs, and the improvement can reach up to 58%. A python implementation of our method is available at https://github.com/Mamba413/AdaDetectGPT.
@@ 1 Introduction
Empirically, we conduct extensive evaluations across multiple datasets and a variety of target language models to demonstrate that AdaDetectGPT consistently outperforms existing logits-based detectors. In white-box settings – where the target LLM to be detected is the same as the source LLM used to compute the logits – AdaDetectGPT achieves improvements over the best alternative in area under the curve (AUC) ranging from 12.5% to 37% over the best existing method. In black-box settings, where the source and target LLMs differ, it can offer gains of up to 58%.
@@ 1.1 Related works
Existing methods for detecting machine-generated text generally fall into three categories: machine learning (ML)-based, statistics-based, and watermarking-based; see Wu et al. (2025) for a recent comprehensive survey. We review the first two categories below and defer the discussion of watermarking-based approaches to Appendix A, since our method, unlike watermarking-based approaches, does not depend on knowledge of the specific hash function or random number generator used during token generation, which are often model-specific and not publicly accessible.
@@
Other statistical measures employed are calculated based on the N-gram distributions (Solaiman et al., 2019; Yang et al., 2024b), the intrinsic dimensionality of text (Tulchinskii et al., 2023), the reward model used in LLMs (Lee et al., 2024) and the maximum mean discrepancy (Zhang et al., 2024; Song et al., 2025). Recent works have extended these methods to more challenging scenarios, such as to handle adversarial attacks (Hu et al., 2023) and black-box settings (Yu et al., 2024; Zeng et al., 2024). Theoretically, Chakraborty et al. (2024) establish a sample complexity bound for detecting machine-generated text.
@@ 2 Preliminaries
We also assume access to a corpus of $n$ human-authored passages $\mathcal{H}=\{\bm{X}^{(i)}\}_{i=1}^{n}\sim p$. This assumption is reasonable, as large corpora of human-written text are readily available online (e.g., Wikipedia). Without loss of generality, we assume that all passages have the same number of tokens $L$, achieved by zero-padding shorter sequences to match the maximum token length. Throughout this paper, we use boldface letters (e.g., $\bm{X}$) to denote passages, and non-boldface letters (e.g., $X$) to denote tokens.
@@ 3 AdaDetectGPT
We present AdaDetectGPT in this section. We begin by discussing the white-box setting in Parts (a)(c): Part (a) introduces the proposed statistical measure; Part (b) discusses our choice of the classification threshold; Part (c) derives a lower bound on the TNR to learn the witness function. Next, in Part (d), we extend our proposal to the black-box setting. Finally, in Part (e), we establish the statistical properties of AdaDetectGPT.
@@
where $w:\mathbb{R}\to\mathbb{R}$ denotes a one-dimensional witness function defined over the space of logits.
@@
where $p_{t}(1)=\mathbb{E}_{X_{<t}\sim p}p(1|X_{<t})$ is determined by the true language distribution $p$. In this case, (5) converges to zero as $q(1)\to 1/2$, regardless of the difference between $q(1)$ and the average of $p_{t}(1)$. As such, there are simple settings in which existing logits-based detectors with an identity witness function will struggle to distinguish between human and machine authored text, independent of the actual distance between $p$ and $q$. However, for any $q(1)\neq 1/2$, there exists a function $w$ that makes (5) equal to $q(1)-L^{-1}\sum_{t=1}^{L}[p_{t}(1)]$, independent of the log ratio (see Appendix B for a formal proof). Thus, whenever $q(1)$ differs from the average $p_{t}(1)$, an appropriate transformation $w$ can reliably detect synthetic humans. We will discuss how to properly learn this witness function in (c).
@@
where the probability is evaluated under the human-generated text distribution $p$. Toward that end, we employ MCLT again to derive a closed-form expression of (7).
@@
Notice that $T_{w}^{(1)}$ is very similar to $T_{w}$ — the only difference lies in the centering term: the conditional expectation is taken with respect to $p_{t}$ instead of $q_{t}$. Under an equal variance condition where the variance terms in the denominator evaluated at $q_{t}$ and $p_{t}$ converge to the same quantity asymptotically, we can invoke the MCLT to prove the asymptotic normality of $T_{w}^{(1)}(\bm{X})$ under the human distribution $p$. As a result, $\textrm{TNR}_{w}$ can be approximated by
@@
Under the equal variance condition and a stochastic dominance condition specified in Appendix D.1, $\textrm{TNR}_{w}$ is asymptotically lower bounded by $\min\{\alpha+\Phi^{\prime}(z_{\alpha})T_{w}^{(2*)},1-\alpha\}$ where $\Phi^{\prime}$ is the derivative of $\Phi$ and $T_{w}^{(2*)}$ denotes a population version of $T_{w}^{(2)}(\bm{X})$, given by
@@
Specifically, we adopt a linear function class $\mathcal{W}=\{w(z)=\phi(z)^{\top}\beta:\|\beta\|_{2}=1\}$ for some bounded $d$-dimensional feature mapping $\phi$. In this case, $\widehat{T}_{w}^{(2)}$ simplifies to $\beta^{\top}\psi/\sqrt{\beta^{\top}\Sigma\beta}$ for some $d$-dimensional vector $\psi$ and $d\times d$ semi-definite positive matrix $\Sigma$ (see Appendix C.2 for the detailed derivation). With some calculations, our estimated regression coefficients $\widehat{\beta}$ can be efficiently obtained by solving the linear system $\Sigma\beta=\psi$, leading to $\widehat{w}=\widehat{\beta}^{\top}\phi$. We set $\phi$ to the B-spline basis function (De Boor, 1978) in our implementation and relegate additional details to Appendix C.2.
(d) Extension to black-box settings. When the target LLM’s logits are unavailable, we propose to employ multiple source LLMs with distributions $q_{1},q_{2},\ldots,q_{m}$, rather than a single source model, to enhance the detection power of AdaDetectGPT. Specifically, for each source distribution $q_{i}$, we apply the procedure in (c) to learn a corresponding witness function $\widehat{w}_{i}$, resulting in the statistic $T_{\widehat{w}_{i}}(\bm{X})$. We next take the maximum over all source models and use $\max_{i}T_{\widehat{w}_{i}}(\bm{X})$ for LLM detection. We empirically investigate this procedure in Section 4.
@@
  • Theorem 2 provides an upper bound on the difference between our classifier’s expected FNR and the target level $\alpha$. This excess FNR depends on two factors: (i) the number of tokens $L$, and (ii) the variance ratio $V_{L}$. As $L$ diverges to infinity and $V_{L}$ stabilizes to $1$, the FNR of our classifier approaches the target level. We remark that assumptions similar to the bounded denominator condition in Theorem 2 are commonly imposed (see e.g., Bolthausen, 1982a; Hall & Heyde, 2014).
@@ 4 Experiments
We conduct numerical experiments in both white-box and black-box settings to illustrate the usefulness of AdaDetectGPT. To save space, some implementation details are provided in Appendix F.
@@
Datasets. We consider three widely-used datasets for comparing different detectors, including SQuAD for Wikipedia-style question answering (Rajpurkar et al., 2016), WritingPrompts for story generation (Fan et al., 2018), XSum for news summarization (Narayan et al., 2018), Yelp for crowd-sourced product reviews (Zhang et al., 2015), and Essay for high school and university-level essays (Verma et al., 2024). Following Bao et al. (2024), we randomly sample 500 human-written paragraphs from each dataset and generate an equal number of machine-authored paragraphs by prompting an LLM with the first 120 tokens of the human-written text and requiring it to complete the text with up to 200 tokens. This is a challenging setting where LLM-generated text is mixed with human writing. To evaluate AdaDetectGPT, we train the witness function $\widehat{w}$ on two of the datasets and calculate the AUC on the remaining testing dataset.
@@
Benchmark methods. We compare the proposed AdaDetectGPT against eight state-of-the-art detectors: Likelihood, Entropy, LogRank (Gehrmann et al., 2019), LogRank ratio (LRR, Su et al., 2023), DetectGPT (Mitchell et al., 2023) and its enhanced variants Normalized Perturbed log Rank (NPR, Su et al., 2023) and Fast-DetectGPT (Bao et al., 2024), DNAGPT (Yang et al., 2024b), Binoculars (Hans et al., 2024), RADAR (Hu et al., 2023), ImBD (Chen et al., 2025a), and BiScope (Guo et al., 2024a). We measure the detection power of each detector using AUC.
@@
Results. We first consider the white-box setting where we employ various source models summarized in Table S6 for text generation, with parameter sizes ranging from 1 to 20 billion. Table 1 reports the AUC scores of various detectors across all combinations of datasets and five source models. It can be seen that AdaDetectGPT achieves the highest AUC across all combinations of datasets and source models, outperforming Fast-DetectGPT – the best baseline method – by 12.5%-37%. We also evaluate AdaDetectGPT on three more advanced open-source LLMs: Qwen2.5 (Bai et al., 2025), Mistral (Jiang et al., 2023), and LLaMA3 (Dubey et al., 2024). As shown in Table S7, AdaDetectGPT delivers uniform improvements over Fast-DetectGPT and maintains competitive performance across five datasets, achieving the best results in most cases. These findings highlight the advantage of using an adaptively learned witness function for classification. In Appendix G.2, we further analyze the computational cost of AdaDetectGPT. Training the witness function typically requires less than one minute across different sample sizes and dimensions, with memory usage below 0.5 GB. I In Appendix G.3, we further conduct a sensitivity analysis to investigate the sensitivity of AdaDetectGPT’s AUC score to various factors that may affect the estimation of the witness function. Our results show that AdaDetectGPT is generally robust to the size of training data, the number of B-spline feature mappings, and the distributional shift between training and test data, consistently maintaining superior performance over baselines.
@@
We next consider the black-box setting, where the task is to detect text generated by three widely used advanced LLMs: GPT-4o (Hurst et al., 2024), Claude-3.5-Haiku (Anthropic, 2024), and Gemini-2.5-Flash (Comanici et al., 2025). In this setting, token-level log-probabilities are not publicly accessible. To implement Fast-DetectGPT, we use google/gemma-2-9b and google/gemma-2-9b-it (Team et al., 2024) as the sampling and scoring models, respectively, to construct the classification statistics described in Section 2. The same setup is also applied when implementing AdaDetectGPT. We further benchmark against existing ML-based classifiers, including GPT-2 detectors based on RoBERTa-base and RoBERTa-large provided by OpenAI. The results are reported in Tables 2 and 3. Overall, Fast-DetectGPT remains the strongest baseline, though it is occasionally outperformed by Binoculars and RADAR. AdaDetectGPT consistently improves upon Fast-DetectGPT on several datasets, with gains on XSum, WritingPrompts, and Essay, reaching as high as 37.8% improvement on Essay when detecting Gemini-2.5. Its performance is also comparable on the Yelp dataset. Finally, we evaluate robustness to adversarial attacks in the black-box setting. As shown in Table S10, AdaDetectGPT demonstrates greater resilience than Fast-DetectGPT, showing reduced sensitivity to adversarially perturbed texts.
Table 2: Performance comparison of GPT-4o and Claude-3.5 across datasets.
Method GPT-4o Claude-3.5 XSum Writing Yelp Essay Avg. XSum Writing Yelp Essay Avg. RoBERTaBase 0.5141 0.5352 0.6029 0.5739 0.5655 0.5206 0.5386 0.5630 0.5593 0.5566 RoBERTaLarge 0.5074 0.5827 0.5027 0.6575 0.5626 0.5462 0.6149 0.5105 0.6063 0.5695 Likelihood 0.5194 0.7661 0.8425 0.7849 0.7282 0.6780 0.7502 0.7134 0.6160 0.6894 Entropy 0.5397 0.7021 0.7291 0.6951 0.6665 0.5935 0.6594 0.6625 0.5742 0.6142 LogRank 0.5123 0.7478 0.8259 0.7786 0.7161 0.5109 0.6653 0.7244 0.7111 0.6529 Binoculars 0.9022 0.9572 0.9840 0.9777 0.9552 0.9012 0.9393 0.9752 0.9603 0.9440 RADAR 0.9580 0.8046 0.8558 0.8394 0.8644 0.9187 0.7264 0.8424 0.9152 0.8507 BiScope 0.8333 0.8733 0.9700 0.9600 0.9092 0.8533 0.8800 0.8800 0.9567 0.8925 Fast-DetectGPT 0.9048 0.9588 0.9847 0.9800 0.9571 0.9019 0.9361 0.9768 0.9608 0.9439 AdaDetectGPT 0.9072 0.9611 0.9832 0.9841 0.9589 0.9176 0.9400 0.9728 0.9610 0.9478 Relative ( ) 2.4288 5.6095 20.4444 4.2454 16.0326 6.0543 0.3405 6.9929
Table 3: Performance of Gemini-2.5 across datasets.
Method XSum Writing Yelp Essay Avg. RoBERTaBase 0.5311 0.5202 0.5624 0.7279 0.5854 RoBERTaLarge 0.6583 0.5888 0.6029 0.8180 0.6845 Likelihood 0.7127 0.7547 0.6566 0.7565 0.7201 Entropy 0.5754 0.5088 0.6023 0.6038 0.5726 LogRank 0.6084 0.5743 0.6896 0.7504 0.6557 Binoculars 0.8500 0.9453 0.9698 0.9908 0.9390 RADAR 0.8184 0.5152 0.6300 0.5891 0.6382 BiScope 0.7633 0.6800 0.7097 0.8167 0.7642 Fast-DetectGPT 0.8404 0.9443 0.9695 0.9914 0.9364 AdaDetectGPT 0.8432 0.9484 0.9644 0.9947 0.9377 Relative ( ) 1.7544 7.4163 37.8238 1.9916
@@
Finally, we employ the same five LLMs from the white-box setting to compare Fast-DetectGPT and AdaDetectGPT in settings. Following Bao et al. (2024), we use GPT-J as the source LLM for detecting each of the remaining four target LLMs. Additionally, we implement a variant of AdaDetectGPT that uses GPT-J for scoring and GPT-Neo for sampling. Due to space constraints, the results are presented Table S11 in Appendix G.5, where AdaDetectGPT uniformly outperforms Fast-DetectGPT in all cases, with improvements of up to 29%.

### v2

@@ Abstract
We study the problem of determining whether a piece of text has been authored by a human or by a large language model (LLM). Existing state of the art logits-based detectors make use of statistics derived from the log-probability of the observed text evaluated using the distribution function of a given source LLM. However, relying solely on log probabilities can be sub-optimal. In response, we introduce AdaDetectGPT – a novel classifier that adaptively learns a witness function from training data to enhance the performance of logits-based detectors. We provide statistical guarantees on its true positive rate, false positive rate, true negative rate and false negative rate. Extensive numerical studies show AdaDetectGPT nearly uniformly improves the state-of-the-art method in various combination of datasets and LLMs, and the improvement can reach up to 37%. A python implementation of our method is available at https://github.com/Mamba413/AdaDetectGPT.
@@ 1 Introduction
Empirically, we conduct extensive evaluations across multiple datasets and a variety of target language models to demonstrate that AdaDetectGPT consistently outperforms existing logits-based detectors. In white-box settings – where the target LLM to be detected is the same as the source LLM used to compute the logits – AdaDetectGPT achieves improvements over the best alternative in area under the curve (AUC) ranging from 12.5% to 37% over the best existing method. In black-box settings, where the source and target LLMs differ, it similarly offer gains of up to 20%.
@@ 1.1 Related works
Existing methods for detecting machine-generated text generally fall into three categories: machine learning (ML)-based, statistics-based, and watermarking-based; see Yang et al. (2024c); Wu et al. (2025) for recent comprehensive surveys. Our method is most closely related to the first two categories and, unlike the third, does not rely on knowledge of the specific hash function or random number generator used during token generation, which are often model-specific and not publicly available. In what follows, we review the first two categories and defer the discussion of watermarking-based approaches to Appendix A.
@@
Other statistical measures employed are calculated based on the N-gram distributions (Solaiman et al., 2019; Yang et al., 2024b), the intrinsic dimensionality of text (Tulchinskii et al., 2023), the reward model used in LLMs (Lee et al., 2024) and the maximum mean discrepancy (Zhang et al., 2024; Song et al., 2025). Recent works have extended these methods to more challenging scenarios, such as to handle adversarial attacks (Hu et al., 2023), machine-revised text (Chen et al., 2025a) and black-box settings (Yu et al., 2024; Zeng et al., 2024). Theoretically, Chakraborty et al. (2024) establish a sample complexity bound for detecting machine-generated text.
@@ 2 Preliminaries
We also assume access to a corpus of $n$ human-authored passages $\mathcal{H}=\{\bm{X}^{(i)}\}_{i=1}^{n}\sim p$. This assumption is reasonable, as large corpora of human-written text are readily available online (e.g., Wikipedia). Without loss of generality, we assume that all passages have the same number of tokens $L$, achieved by zero-padding shorter sequences to match the maximum token length. Throughout this paper, we use boldface letters (e.g., $\bm{X}$) to denote passages, and non-boldface letters (e.g., $X$) to denote individual tokens.
@@ 3 AdaDetectGPT
We present AdaDetectGPT in this section. We begin by discussing the white-box setting in Parts (a)(c): Part (a) introduces the proposed statistical measure; Part (b) discusses our choice of the classification threshold for FNR control; Part (c) derives a lower bound on the TNR to learn the witness function. Next, in Part (d), we extend our proposal to the black-box setting. Finally, in Part (e), we establish the statistical properties of AdaDetectGPT.
@@
where $w:\mathbb{R}\to\mathbb{R}$ denotes a one-dimensional witness function defined over the space of log-probabilities.
@@
where $p_{t}(1)=\mathbb{E}_{X_{<t}\sim p}p(1|X_{<t})$ is determined by the true language distribution pp. In this case, (5) converges to zero as $q(1)\to 1/2$, regardless of the difference between $q(1)$ and the average of $p_{t}(1)$. As such, there are simple settings in which existing logits-based detectors with an identity witness function will struggle to distinguish between human and machine authored text, independent of the actual distance between $p$ and $q$. However, for any $q(1)\neq 1/2$, there exists a function $w$ that makes (5) equal to $q(1)-L^{-1}\sum_{t=1}^{L}[p_{t}(1)]$, independent of the log ratio (see Appendix B for a formal proof). Thus, whenever $q(1)$ differs from the average $p_{t}(1)$, an appropriate transformation $w$ can reliably detect synthetic humans.
We will discuss how to properly learn the witness function below. Since this is inherently a classification problem, it is natural to learn a witness function that maximizes the AUC of the resulting detector. Equivalently, this amounts to finding a witness function that maximizes the TNR for each fixed FNR. In (b), we discuss how to select the classification threshold to control the FNR at a specified level. Given this threshold, we then derive a lower bound on the TNR in (c) to be maximized for learning the witness function.
@@
where the probability is evaluated under the human-generated text distribution pp. Toward that end, we employ MCLT again to derive a closed-form expression of (7).
@@
Notice that $T_{w}^{(1)}$ is very similar to $T_{w}$ — the only difference lies in the centering term: the conditional expectation is taken with respect to $p_{t}$ instead of $q_{t}$. Under an equal variance condition where the variance terms in the denominator evaluated at $q_{t}$ and $p_{t}$ converge to the same quantity asymptotically, we can invoke the MCLT to prove the asymptotic normality of $T_{w}^{(1)}(\bm{X})$ under the human distribution pp. As a result, $\textrm{TNR}_{w}$ can be approximated by
@@
Under the equal variance condition specified in Appendix D, $\textrm{TNR}_{w}$ is asymptotically lower bounded by $\min\{\alpha+\Phi^{\prime}(z_{\alpha})T_{w}^{(2*)},1-\alpha\}$ where $\Phi^{\prime}$ is the derivative of $\Phi$ and $T_{w}^{(2*)}$ denotes a population version of $T_{w}^{(2)}(\bm{X})$, given by
@@
Specifically, we adopt a linear function class $\mathcal{W}=\{w(z)=\phi(z)^{\top}\beta:\|\beta\|_{2}=1\}$ for some bounded $d$-dimensional feature mapping $\phi$. In this case, $\widehat{T}_{w}^{(2)}$ simplifies to $\beta^{\top}\psi/\sqrt{\beta^{\top}\Sigma\beta}$ for some $d$-dimensional vector $\psi$ and $d\times d$ semi-definite positive matrix $\Sigma$ (see Appendix C for the detailed derivation). With some calculations, our estimated regression coefficients $\widehat{\beta}$ can be efficiently obtained by solving the linear system $\Sigma\beta=\psi$, leading to $\widehat{w}=\widehat{\beta}^{\top}\phi$. We set $\phi$ to the B-spline basis function (De Boor, 1978) in our implementation and relegate additional details to Appendix C.
(d) Extension to black-box settings. When the target LLM’s logits are unavailable, we employ an open-source LLM with a distribution similar to that of the target model to construct the statistical measure in (4). The witness function is then learned in the same manner as described in (c). We empirically evaluate this approach in Section 4.
@@
  • Theorem 2 provides an upper bound on the difference between our classifier’s expected FNR and the target level $\alpha$. This excess FNR depends on two factors: (i) the number of tokens $L$, and (ii) the variance ratio $V_{L}$. As $L$ diverges to infinity and $V_{L}$ stabilizes to 11, the FNR of our classifier approaches the target level. We remark that assumptions similar to the bounded denominator condition in Theorem 2 are commonly imposed (see e.g., Bolthausen, 1982; Hall & Heyde, 2014).
@@ 4 Experiments
We conduct numerical experiments in both white-box and black-box settings to illustrate the usefulness of AdaDetectGPT. To save space, some implementation details are provided in Appendix C.
@@
Datasets. We consider five widely-used datasets for comparing different detectors, including SQuAD for Wikipedia-style question answering (Rajpurkar et al., 2016), WritingPrompts for story generation (Fan et al., 2018), XSum for news summarization (Narayan et al., 2018), Yelp for crowd-sourced product reviews (Zhang et al., 2015), and Essay for high school and university-level essays (Verma et al., 2024). Following Bao et al. (2024), we randomly sample 500 human-written paragraphs from each dataset and generate an equal number of machine-authored paragraphs by prompting an LLM with the first 120 tokens of the human-written text and requiring it to complete the text with up to 200 tokens. This is a challenging setting where LLM-generated text is mixed with human writing. To evaluate AdaDetectGPT, we compute the AUC on each of the five datasets, with its witness function $\widehat{w}$ trained on two randomly selected datasets that differ from the test dataset.
@@
Benchmark methods. In white-box settings, we compare the proposed AdaDetectGPT against eight state-of-the-art detectors: Likelihood, Entropy, LogRank (Gehrmann et al., 2019), LogRank Ratio (LRR, Su et al., 2023), DetectGPT (Mitchell et al., 2023) and its variants Normalized Perturbed log Rank (NPR, Su et al., 2023), Fast-DetectGPT (Bao et al., 2024), DNAGPT (Yang et al., 2024b). In black-box settings, we further compare against RoBERTaBase and RoBERTaLarge (Solaiman et al., 2019), Binoculars (Hans et al., 2024), RADAR (Hu et al., 2023), and BiScope (Guo et al., 2024a), but omit DetectGPT, NPR and DNAGPT due to their high computational cost. This yields ten baseline algorithms. We measure the detection power of each detector using AUC.
@@
White-box results. We first consider the white-box setting where we employ various source models summarized in Table S4 for text generation, with parameter sizes ranging from 1 to 20 billion. Table 1 reports the AUC scores of various detectors across all combinations of datasets and five source models. It can be seen that AdaDetectGPT achieves the highest AUC across all combinations of datasets and source models, outperforming Fast-DetectGPT – the best baseline method – by 12.5%-37%. We also evaluate AdaDetectGPT on three more advanced open-source LLMs: Qwen2.5 (Bai et al., 2025), Mistral (Jiang et al., 2023), and LLaMA3 (Grattafiori et al., 2024). As shown in Table S7, AdaDetectGPT delivers consistent improvements over Fast-DetectGPT and maintains competitive performance across five datasets, achieving the best results in most cases. These findings highlight the advantage of using an adaptively learned witness function for classification.
In Appendix F.3, we analyze the computational cost of AdaDetectGPT. Training the witness function typically requires less than one minute across different sample sizes and dimensions, with memory usage below 0.5 GB. In Appendix F.4, we further conduct a sensitivity analysis to investigate the sensitivity of AdaDetectGPT’s AUC score to various factors that may affect the estimation of the witness function. Our results show that AdaDetectGPT is generally robust to the size of training data, the number of B-spline features, and the distributional shift between training and test data, consistently maintaining superior performance over baselines.
@@
Black-box results. We next consider the black-box setting, where the task is to detect text generated by three widely used advanced LLMs: GPT-4o (Hurst et al., 2024), Claude-3.5-Haiku (Anthropic, 2024), and Gemini-2.5-Flash (Comanici et al., 2025). In this setting, token-level log-probabilities are not publicly accessible. To implement Fast-DetectGPT and AdaDetectGPT, we use google/gemma-2-9b and google/gemma-2-9b-it (Team et al., 2024) as the sampling and scoring models, respectively, to construct the classification statistics described in Section 2. The results are reported in Tables 2 and S8. Overall, Fast-DetectGPT remains the strongest baseline, although it occasionally underperforms Binoculars or RADAR. Nonetheless, AdaDetectGPT consistently improves upon Fast-DetectGPT across various datasets, with gains on Essay reaching up to 37.8%.
@@
Additionally, we evaluate AdaDetectGPT’s robustness to two adversarial attacks, paraphrasing and decoherence, in the black-box setting. As shown in Table 3, AdaDetectGPT demonstrates greater resilience than Fast-DetectGPT to adversarially perturbed texts. The improvement reaches up to 10% for paraphrasing and up to 85% for decoherence. Finally, we employ the same five LLMs from Table 1 to compare Fast-DetectGPT and AdaDetectGPT in black-box settings. Following Bao et al. (2024), we use GPT-J as the source LLM for detecting each of the remaining four target LLMs. Due to space constraints, the results are presented in Table S11 of Appendix F.5, where AdaDetectGPT uniformly outperforms Fast-DetectGPT in all cases, with improvements of up to 29%.
Table 2: AUC scores of various detectors to detect text generated by GPT-4o and Claude-3.5 across datasets.
Method GPT-4o Claude-3.5 XSum Writing Yelp Essay Avg. XSum Writing Yelp Essay Avg. RoBERTaBase 0.5141 0.5352 0.6029 0.5739 0.5655 0.5206 0.5386 0.5630 0.5593 0.5566 RoBERTaLarge 0.5074 0.5827 0.5027 0.6575 0.5626 0.5462 0.6149 0.5105 0.6063 0.5695 Likelihood 0.5194 0.7661 0.8425 0.7849 0.7282 0.6780 0.7502 0.7134 0.6160 0.6894 Entropy 0.5397 0.7021 0.7291 0.6951 0.6665 0.5935 0.6594 0.6625 0.5742 0.6142 LogRank 0.5123 0.7478 0.8259 0.7786 0.7161 0.5109 0.6653 0.7244 0.7111 0.6529 LRR 0.5116 0.6099 0.6828 0.6930 0.6243 0.5268 0.5560 0.5527 0.6535 0.5723 Binoculars 0.9022 0.9572 0.9840 0.9777 0.9552 0.9012 0.9393 0.9752 0.9603 0.9440 RADAR 0.9580 0.8046 0.8558 0.8394 0.8644 0.9187 0.7264 0.8424 0.9152 0.8507 BiScope 0.8333 0.8733 0.9700 0.9600 0.9092 0.8533 0.8800 0.8800 0.9567 0.8925 Fast-DetectGPT 0.9048 0.9588 0.9847 0.9800 0.9571 0.9019 0.9361 0.9768 0.9608 0.9439 AdaDetectGPT 0.9072 0.9611 0.9832 0.9841 0.9589 0.9176 0.9400 0.9728 0.9610 0.9478 Relative ( ) 2.4288 5.6095 20.4444 4.2454 16.0326 6.0543 0.3405 6.9929
Table 3: Detection of LLM-generated text under two adversarial attacks in black-box settings.
Paraphrasing Decoherence
DetectGPT Xsum Writing PubMed Avg. Xsum Writing PubMed Avg.